OpenStudy (anonymous):

Whats the first step that i should do with this integral? I tried to solve it by parts but i dont know what U should be.

OpenStudy (anonymous):

\[\int\limits (x ^{2}+1)COS(x ^{3}+3x)\]

OpenStudy (anonymous):

first decide which one you wanna make as variable, for example x2+1. then you solve the cos (x3+3x) integral

OpenStudy (anonymous):

after finish, you make the cos (x3+3x) as variable, then finish x2+1 integral

OpenStudy (anonymous):

u = x³ + 3x --> du = ...?

OpenStudy (anonymous):

its differential. du=3x2+3

OpenStudy (anonymous):

@Chef86 Hello...0!

OpenStudy (anonymous):

@Chef86 Why don't you show up on your post instead of messaging?

OpenStudy (anonymous):

\[dU = \frac{ SIN(x ^{3}+3x) }{ 3x ^{2}+3 }\]

OpenStudy (anonymous):

dang i meant T

OpenStudy (anonymous):

No, u = x³ + 3x --> du = ...? ( no sin, cos yet)

OpenStudy (anonymous):

du = 2x^2 + 3

OpenStudy (anonymous):

its 3x^2 +3

OpenStudy (anonymous):

whoops i meant 3

OpenStudy (anonymous):

3(x^2 +1)

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 } du\]

OpenStudy (anonymous):

x^2 + 1 reminds me of the inverse of the tangent

OpenStudy (anonymous):

meeen, its my first year calculus university, I've alrready forget >.<

OpenStudy (anonymous):

NO, don't go over complicated !!! u = x³ + 3x -->du/ 3 = ( x² + 1 ) dx Just plug u and du in!

OpenStudy (anonymous):

It's plain U-substitution!

OpenStudy (anonymous):

i was trying to do it by parts before which is why i got confused by this

OpenStudy (anonymous):

and i just put the sin in and its done basicly

OpenStudy (anonymous):

sin(U) + c

OpenStudy (anonymous):

So 1/3 ∫ cos u du = (1/3) sin u+ C = (1/3) sin ( x³ + 3x) + C

OpenStudy (anonymous):

thats what i got thks

OpenStudy (anonymous):

Then why didn't you post it, instead of going around by Part, then U !!!

OpenStudy (anonymous):

Im sorry that i even asked, i didnt mean to get you so angry.